OpenStudy (anonymous):

Please Help! What's the difference between the Symmetric and Reflexive Properties?

OpenStudy (anonymous):

Reflexive Property states that anything equals itself. a = a

OpenStudy (anonymous):

says that if a number is equal to a second number, then the second number is equal to the first number. If a=b, then b=a

OpenStudy (anonymous):

Fang02

OpenStudy (anonymous):

Okay, I understand now. Thanks! (I was eating lunch, so that's why I was gone.) Could you please confirm the following? They want me to complete the two-column proof. Given: 11x - 6y = -1; x = 8 Prove: 89/6 = y After the Given they state the steps. They want me to state what property each step is. 88 - 6y = -1 This is Substitution I think. -6 = -89 I'm not sure about this one. y = 86/6 Division Property 89/6 = y Symmetric Am I correct?